Can someone help me solve this problem?
Steve traveled 150 miles at a certain speed. Had he gone 20mph faster, the trip would have taken 2 hrs less. Find the speed of his vehicle.

Answer by RAY100(1637)

Velocity = distance / Time ,,,,or t= d/v
.
T1 -2 = T2
.
150/v -2 = 150 /(v+20)
.
mult thru by v(v+20)
.
150(v+20) -2(v)(v+20) = 150(v)
.
150v +3000 -2v^2 -40v =150v
.
-2v^2 -40v +3000 = 0
.
v^2 +20v -1500 = 0
.
(v+50)(v-30) = 0
.
v= 30, -50 ( not realistic)
.
check
150/30=5,,,150/(30+20) =3,,,,ok